### v1

# Corrigendum for Han’s Corrigendum

Abstract
S.-E. Han’s paper, ”Remarks on Pseudocovering Spaces in a Digital Topological Setting: A Corrigendum”, is meant to address errors in previous papers. However, this paper is also marked by errors in its mathematics, as well as improprieties in its citations. We address these flaws in the current work.
Key words and phrases: digital topology, digital image, covering map
MSC: 54B20, 54C35
1 Introduction
S.-E. Han’s paper [19] claims in its title to be a Corrigendum for his papers [16, 18], but is itself in need of a Corrigendum. These papers are concerned with attempted variants on covering spaces in digital topology, among which are those that turn out to coincide with digital covering spaces. The current paper discusses flaws in Han’s Corrigendum.
The notion of a covering map has been adapted from classical algebraic topology to digital topology, where it is an important tool for computing digital versions of fundamental groups for binary digital images. [6]
We show that Han’s “Corrigendum” remains flawed in multiple ways:
  • One of the “proofs” in [19] is incorrect, although its assertion is correct; we discuss the error and give a correct proof of the assertion.
  • Much of [19] is unoriginal, with original versions not credited.
  • Han doesn’t even cite his own papers correctly.
2 Preliminaries
2.1 Digital images
Material in this section is largely quoted or paraphrased from [6].
We use $\mathbb{N}$} for the set of natural numbers, $\mathbb{Z}$} for the set of integers, and #X\#X for the number of distinct members of $X$. Let ${\mathbb{N}}^{*}={\mathbb{N}}\cup\{0\}$.
We typically denote a (binary) digital image as $(X,\kappa)$, where $\emptyset\neq X\subset{\mathbb{Z}}^{n}$ for some $n\in{\mathbb{N}}$ and $\kappa$ represents an adjacency relation of pairs of points in $X$. Thus, $(X,\kappa)$ is a graph, in which members of $X$ may be thought of as black points, and members of ${\mathbb{Z}}^{n}\setminus X$ as white points, typically of a picture of some “real world” object or scene.
Let $u,n\in{\mathbb{N}}$, $1\leq u\leq n$. Han’s papers use “$k$-adjacency” sometimes to mean an arbitrary adjacency, sometimes as an abbreviation for what he calls “$k(u,n)$-adjacency,” where the digital image $(X,k)$ satisfies $X\subset{\mathbb{Z}}^{n}$ and $x=(x_{1},\ldots,x_{n}),~y=(y_{1},\ldots,y_{n})\in X$ are $k(u,n)$-adjacent if and only if
  • $x\neq y$, and
  • for at most $u$ indices $i$, $\mid x_{i}-y_{i}\mid=1$, and
  • for all indices $j$ such that $\mid x_{j}-y_{j}\mid\neq 1$, we have $x_{j}=y_{j}$.
Generally, $n$ is either explicitly given (e.g., $n=2$) or is understood as an abstract positive integer. Therefore, other authors refer to this adjacency as $c_{u}$-adjacency. We will prefer the latter notation in the current paper. The $c_{u}$ adjacencies are the adjacencies most used in digital topology, especially $c_{1}$ and $c_{n}$.
In low dimensions, it is also common to denote a $c_{u}$ adjacency by the number of points that can have this adjacency with a given point in ${\mathbb{Z}}^{n}$. E.g.,
  • For subsets of ${\mathbb{Z}}^{1}$, $c_{1}$-adjacency is 2-adjacency.
  • For subsets of ${\mathbb{Z}}^{2}$, $c_{1}$-adjacency is 4-adjacency and $c_{2}$-adjacency is 8-adjacency.
  • For subsets of ${\mathbb{Z}}^{3}$, $c_{1}$-adjacency is 6-adjacency, $c_{2}$-adjacency is 18-adjacency, and $c_{3}$-adjacency is 26-adjacency.
We use the notations $y\leftrightarrow_{\kappa}x$, or, when the adjacency $\kappa$ can be assumed, $y\leftrightarrow x$, to mean $x$ and $y$ are $\kappa$-adjacent. The notations $y\leftrightarroweq_{\kappa}x$, or, when $\kappa$ can be assumed, $y\leftrightarroweq x$, mean either $y=x$ or $y\leftrightarrow_{\kappa}x$. For $x\in X$, let
$N(X,x,\kappa)=\{\,y\in X\mid x\leftrightarrow_{\kappa}y\,\}.$
When the image $(X,\kappa)$ under discussion is clear, we will use the notations $N(x)$ or $N_{\kappa}(x)$ as follows.
$N(x)=\{\,y\in X\mid y\leftrightarroweq_{\kappa}x\,\}=N(X,x,\kappa)\cup\{x\}.$
A sequence $P=\{y_{i}\}_{i=0}^{m}$ in a digital image $(X,\kappa)$ is a $\kappa$-path from $a\in X$ to $b\in X$ if $a=y_{0}$, $b=y_{m}$, and $y_{i}\leftrightarroweq_{\kappa}y_{i+1}$ for $0\leq i<m$.
$X$ is $\kappa$-connected [24], or connected when $\kappa$ is understood, if for every pair of points $a,b\in X$ there exists a $\kappa$-path in $X$ from $a$ to $b$.
A (digital) $\kappa$-closed curve is a path $S=\{s_{i}\}_{i=0}^{m}$ such that $s_{0}=s_{m}$, and 0<|ij|<m0<|i-j|<m implies $s_{i}\neq s_{j}$. If, also, $0\leq i<m$ implies
$N(S,x_{i},\kappa)=\{x_{(i-1)\mod n},~x_{(i+1)\mod m}\}$
then $S$ is a (digital) $\kappa$-simple closed curve.
2.2 Digitally continuous functions
Material in this section is largely quoted or paraphrased from [6].
Digital continuity is defined to preserve connectedness, as at Definition 2.1 below. By using adjacency as our standard of “closeness,” we get Theorem 2.2 below.
Definition 2.1.
[2] (generalizing a definition of [24]) Let $(X,\kappa)$ and $(Y,\lambda)$ be digital images. A function $f:X\rightarrow Y$ is $(\kappa,\lambda)$-continuous if for every $\kappa$-connected $A\subset X$ we have that $f(A)$ is a $\lambda$-connected subset of $Y$.
If either of $X$ or $Y$ is a subset of the other, we use the abbreviation $\kappa$-continuous for $(\kappa,\kappa)$-continuous.
When the adjacency relations are understood, we will simply say that $f$ is continuous. Continuity can be expressed in terms of adjacency of points:
Theorem $2.$
[24, 2] A function $f:X\to Y$ is continuous if and only if $x\leftrightarrow x^{\prime}$ in $X$ implies $f(x)\leftrightarroweq f(x^{\prime})$.
Han’s papers generally use the equivalent formulation that $f$ is continuous if and only for every $x\in X$, $f(N_{\kappa}(x))\subset N_{\lambda}(f(x))$.
See also [8, 9], where similar notions are referred to as immersions, gradually varied operators, and gradually varied mappings.
A digital isomorphism (called homeomorphism in [1]) is a $(\kappa,\lambda)$-continuous surjection $f:X\to Y$ such that $f^{-1}:Y\to X$ is $(\lambda,\kappa)$-continuous.
The literature uses path polymorphically: a $(c_{1},\kappa)$-continuous function $f:[0,m]_{{\mathbb{Z}}}\to X$ is a $\kappa$-path if $f([0,m]_{{\mathbb{Z}}})$ is a $\kappa$-path as described above from $f(0)$ to $f(m)$.
2.3 Han’s variants on local isomorphisms
Material in this section is largely quoted or paraphrased from [6].
The following Definition 2.3 of a digital covering map is formally simpler but equivalent to the definition Han introduced in [13]. The equivalence was shown in [3] and at about the same time in [14]; it may have been derived independently by the authors of these two papers. However, in [19], Definition 2.3 is attributed to three of Han’s papers without acknowledging its origin in [3, 14].
Definition 2.3.
[3] Let $p:(E,\kappa)\to(B,\lambda)$ be a continuous surjection of digital images. The map $p$ is a $(\kappa,\lambda)$-covering (map) if and only if
  1. 1.
    for every $b\in B$, there is an index set $M$ such that
    $p^{-1}(N_{\lambda}(b))=\bigcup_{i\in M}N_{\kappa}(e_{i}),\mbox{ where }e_{i}\in p^{-1}(b);$
  2. 2.
    if $\in M, iji\neq j$, then $N_{\kappa}(e_{i})\cap N_{\kappa}(e_{j})=\emptyset$; and
  3. 3.
    $p|_{N_{\kappa}(e_{i})}:N_{\kappa}(e_{i})\to N_{\lambda}(b)$ is a $(\kappa,\lambda)$-isomorphism.
We find the following definition in Han’s paper [17] (not in [12] despite the claims to the contrary in [17, 18]).
Definition 2.4.
A digitally continuous map $h:(X,\kappa)\to(Y,\lambda)$ is a pseudo-local (PL) isomorphism if for every $x\in X$, $h(N_{\kappa}(x))\subset Y$ is $\lambda$-isomorphic to $N_{\lambda}(h(x))\subset Y$.
In his paper [12], Han gives the following.
Definition 2.5.
A digitally continuous map $h:(X,\kappa)\to(Y,\lambda)$ is a local homeomorphism [in more recent terminology, a local isomorphism] if for all $x\in X$, $h|_{N_{\kappa}(x)}$ is a $(\kappa,\lambda)$-homeomorphism $(\kappa,\lambda)$-isomorphism] onto $N_{\lambda}(h(x))$.
The following was noted in [6].
Proposition 2.6.
Let $h:(X,\kappa)\to(Y,\lambda)$ be a digitally continuous map. If $h$ is a local isomorphism then $h$ is a PL isomorphism.
Theorem 2.7.
([23], correcting an error of [12]) Let $f:(X,\kappa)\to(Y,\lambda)$ be a continuous surjection. Then $f$ is a digital covering map if and only if $f$ is a local isomorphism.
We will also discuss the following notion.
Definition 2.8.
[16] A function $h:(X,\kappa)\to(Y,\lambda)$ is a weakly local (WL) isomorphism if for all $x\in X$, $h|_{N(x,1)}$ is an isomorphism onto $h(N(x,1))$.
3 Han’s pseudo-covering maps in  [18]
Han defines a digital pseudo-covering as follows.
Definition 3.1.
[16] Let $p:(E,\kappa)\to(B,\lambda)$ be a surjection such that for every $b\in B$,
  1. 1.
    there is an index set $M$ such that $p^{-1}(N_{\kappa}(b,1))=\bigcup_{i\in M}N_{\kappa}(e_{i},1)$, where $e_{i}\in p^{-1}(b)$;
  2. 2.
    if $i,j\in M$ and $i\neq j$, then $N_{\kappa}(e_{i},1)\cap N_{\kappa}(e_{j},1)=\emptyset$; and
  3. 3.
    $p|_{N_{\kappa}(e_{i},1)}:N_{\kappa}(e_{i},1)\to N_{\lambda}(b,1)$ is a WL-isomorphism for all $i\in M$.
Then $p$ is a pseudo-covering map.
However, A. Pakdaman shows in [21] that Han’s definition does not effectively give us a new object of study. In particular, Pakdaman shows the following.
Theorem 3.2.
[21] A digital pseudo-covering map as defined in Definition 3.1 is in fact a digital covering map.
Definition 3.3.
[13] Let $p:(E,\kappa)\to(B,\lambda)$ be $(\kappa,\lambda)$-continuous. Let $f:[0,m]_{{\mathbb{Z}}}\to B$ be $(c_{1},\lambda)$-continuous. A $(c_{1},\kappa)$-continuous function $\tilde{f}:[0,m]_{{\mathbb{Z}}}\to E$ such that $p\circ\tilde{f}=f$ is a (digital) path lifting of $f$. If for every $b_{0}\in B$, every $e_{0}\in p^{-1}(b_{0})$, and every path $f$ such that $f(0)=b_{0}$, there is a unique lifting $\tilde{f}$ such that $\tilde{f}(0)=e_{0}$, then $p$ has the unique path lifting property.
Han’s variants of covering maps don’t all give us new notions, as shown by the following expansion of Theorem 3.2.
Theorem 3.4.
[6] Let $p:(X,\kappa)\to(Y,\lambda)$ be a continuous surjection. Then the following are equivalent.
  • $p$ is a digital covering map.
  • $p$ is a local isomorphism.
  • $p$ is a pseudo-covering in the sense of Definition 3.1.
  • $p$ is a WL-isomorphism with the unique path lifting property.
In order to obtain a notion of a digital pseudo-cover that is not equivalent to a digital cover, Pakdaman gave the following definition.
Definition 3.5.
[21] Let $p:(E,\kappa)\to(B,\lambda)$ be a surjection of digital images. Suppose for all $b\in B$ we have the following.
  1. 1.
    for some index set $M$, $\bigcup_{i\in M}N_{\kappa}(e_{i},1)\subset p^{-1}(N_{\lambda}(b,1))$ where $e_{i}\in p^{-1}(b)$;
  2. 2.
    if $i,j\in M$ and $i\neq j$ then $N_{\kappa}(e_{i},1)\cap N_{\kappa}(e_{i},1)=\emptyset$; and
  3. 3.
    for all $i\in M$, $p|{N_{\kappa}(e_{i},1)}:N_{\kappa}(e_{i},1)\to p(N_{\kappa}(e_{i},1))$ is a $(\kappa,\lambda)$-isomorphism.
Then $p$ is a $(\kappa,\lambda)$-pseudocovering map.
Pakdaman [21] showed that Definition 3.5, unlike Definition 3.1, gives us a notion of a digital pseudo-cover that is not a digital cover. He did this by showing that, unlike a digital cover, the version of digital pseudo-cover in Definition 3.5 need not have the unique lifting property.
Having recognized that his previous definition of a digital pseudo-cover, stated here as Definition 3.1, failed to achieve his goal, Han then stated in [19] the same alternate definition of a digital pseudo-cover that Pakdaman gave (above, Definition 3.5), without attributing it to Pakdaman (see Han’s Definition 4.1 in [19]). Since [19] does acknowledge the existence of Pakdaman’s [21], this serves as an example of, at best, careless research by Han, and perhaps a breach of ethics.
4 Other flaws in [19]
In this section, we take $p:{\mathbb{N}}^{*}\to C_{n}$, where $C_{n}$ is a digital simple closed curve of $n$ points $\{c_{0},c_{1},\ldots,c_{n-1}\}$ indexed circularly, and $p$ is defined by
$p(z)=c_{z\!\!\mod n}.$
In Remark 3.10(1) of [19], Han claims
Assertion 4.1.
There exists $b\in C_{n}$ such that
$\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i})\neq p^{-1}(N(b)),~~~\mbox{ where }~~~p^{-1}(b)=\{e_{i}\mid i\in{\mathbb{N}}^{*}\}.$
Proposition 4.2.
Assertion 4.1 is false.
Proof.
Consider the following cases.
  • Suppose $b=c_{0}$. We have
    $p^{-1}(N(c_{0}))=\{0,1\}\cup\bigcup_{i=1}^{\infty}[i\cdot n-1,i\cdot n+1]_{{\mathbb{Z}}}=\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i}),$
    where $e_{i}=i\cdot n$ and $p^{-1}(0)=\{e_{i}\mid i\in{\mathbb{N}}^{*}\}$.
  • Suppose $b=c_{j}$ such that $j\neq 0\mod n$. We have
    $p^{-1}(N(c_{j}))=\bigcup_{i=0}^{\infty}[i\cdot j-1,i\cdot n+1]_{{\mathbb{Z}}}=\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i}),$
    where $i\mod n=j$ and $p^{-1}(c_{j})=\{e_{i}\mid i\in{\mathbb{N}}^{*}\}$.
Therefore we have $p^{-1}(N(b))=\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i})$ in both cases. This completes the proof. ∎
Han claims that the following Corollary 4.3 (stated in[19] as Corollary 3.11) follows from his Remark 3.10. However, since we see in Proposition 4.2 that Han’s Remark 3.10 is false, Corollary 4.3 remains in need of a proof, which we give below.
Corollary 4.3.
A pseudo-covering map in the sense of Definition 3.1 is a WL-isomorphic surjection, but the converse does not obtain.
Proof.
By Theorem 3.4, a pseudo-covering map in the sense of Definition 3.1 is a WL-isomorphic surjection.
The map $p$ assumed for this section is easily seen to be a WL-isomorphic surjection. However, as in the proof of Proposition 3.2 of [21]$p$ lacks the unique path lifting property: Given the path $P=\{s_{0},s_{n-1}\}$ in $C_{n}$, there is no lift of $P$ to a path in $N^{*}$ from $0\in p^{-1}(s_{0})$ to any member of $p^{-1}(s_{n-1})$. It follows from Theorem 3.4 that $p$ is not a pseudo-covering map in the sense of Definition 3.1. ∎
5 Further remarks
S.-E. Han’s paper [19] is shown in the current work to have a major mathematical error, trivialities (having introduced variants on the notion of a covering map that don’t really vary), content that is unoriginal but presented as if original, and citations that are incorrect or inappropriate. These are features of many of Han’s papers, examples of which, in addition to those discussed above, include:
Han paper Flaws discussed in
[11] [4]
[13] [3]
[15] [7]
[20] [5]
Perhaps the outstanding example of Han’s unethical conduct is in his paper [10], which is difficult to find as its vehicle, the Journal of Computer &\& Communications Research, is not an international journal; rather, this “journal” is described by Han as a local publication of Honam University, at which Han was employed when the paper was written. This paper is largely plagiarized from [1, 2] and has often been cited by Han after he pledged not to do so – documentation available upon request.
Authors are responsible for their writings, but referees who approve publication of mathematical trash also bear some responsibility. The point is not that a Han submission should be automatically rejected - he deserves credit for useful ideas such as the introduction of covering maps and the shortest-path metric into digital topology. Rather, his submissions should be reviewed with the greatest care, by reviewers with good knowledge of the relevant literature, who are not afraid to recommend either rejection or major revision.